# 3DGAA: Realistic and Robust 3D Gaussian-based Adversarial Attack for Autonomous Driving

### v1

@@ Abstract
Camera-based object detection systems play a vital role in autonomous driving, yet they remain vulnerable to adversarial threats in real-world environments. While existing 2D and 3D physical attacks typically optimize texture, they often struggle to balance physical realism and attack robustness. In this work, we propose 3D Gaussian-based Adversarial Attack (3DGAA), a novel adversarial object generation framework that leverages the full 14-dimensional parameterization of 3D Gaussian Splatting (3DGS) to jointly optimize geometry and appearance in physically realizable ways. Unlike prior works that rely on patches or texture, 3DGAA jointly perturbs both geometric attributes (shape, scale, rotation) and appearance attributes (color, opacity) to produce physically realistic and transferable adversarial objects. We further introduce a physical filtering module to preserve geometric fidelity, and a physical augmentation module to simulate complex physical scenarios, thus enhancing attack generalization under real-world conditions. We evaluate 3DGAA on both virtual benchmarks and physical-world setups using miniature vehicle models. Experimental results show that 3DGAA achieves to reduce the detection mAP from 87.21% to 7.38%, significantly outperforming existing 3D physical attacks. Moreover, our method maintains high transferability across different physical conditions, demonstrating a new state-of-the-art in physically realizable adversarial attacks. These results validate 3DGAA as a practical attack framework for evaluating the safety of perception systems in autonomous driving.
@@ 1 Introduction
Camera-based object detection systems serve as a cornerstone of autonomous driving [2, 1, 30], enabling essential tasks such as obstacle avoidance and lane navigation. However, increasing research has shown that these models are highly vulnerable to adversarial examples–subtle, structured perturbations that induce misclassifications even under physical world conditions [4, 8]. This raises serious concerns about the reliability of perception modules in safety-critical scenarios, especially in the presence of adversarial attacks that can actively exploit their weaknesses. Thus, systematically exploring such vulnerabilities [3, 6, 31, 37, 12, 35, 33, 32, 26, 27, 13] is essential to understand and improve the real world safety of autonomous driving systems.
@@
Existing physical adversarial attacks typically fall into two categories: 2D patch-based methods and 3D texture-level attacks. 2D methods typically apply adversarial patches onto objects [3, 6, 31, 12], achieving high physical realism through minimal appearance changes, but often suffer from poor robustness across viewpoints or physical conditions (Challenge 1). In contrast, recent 3D attacks [37, 35, 33, 32, 26, 27] achieve stronger and more consistent multi-view adversarial performance by directly manipulating surface textures. However, they often rely solely on appearance perturbations, which may result in visual artifacts or unrealistic surface distortions. This limits their applicability in safety-critical settings, where high physical realism is essential for undetectable deployment (Challenge 2). This reveals an inherent trade-off between physical realism and adversarial robustness, posing a key challenge for deploying such attacks in real-world scenarios where both are critical (Challenge 3).
@@
To address these challenges, we introduce 3DGAA, an adversarial attack framework that leverages 3D Gaussian Splatting (3DGS)–a differentiable, compact 3D representation originally developed for photorealistic rendering [16]. Compared to mesh-based or point-based representations, 3DGS naturally encodes both geometry and appearance within a unified 14-dimensional space, allowing fine-grained and jointly optimized perturbations that are crucial for the generation of physically realistic adversarial objects [11, 29, 28].
@@
3DGAA comprises three key modules that jointly address the trade-off between physical realism and attack robustness: First, a Physically-Constrained Adversarial Optimization stage jointly perturbs geometry and appearance attributes to degrade detector confidence across diverse viewpoints and camera distances. This overcomes the limited perturbation scope of 2D methods and addresses Challenge 1. Second, we propose a Physical Filtering Module that enhances geometric fidelity by removing topological outliers and denoising structural artifacts. This module enforces surface-level consistency and mitigates unrealistic deformations commonly seen in texture-only adversarial models, thereby preserving the visual plausibility of the 3D object from multiple perspectives, further enhancing the physical reality, tackling Challenge 2. Furthermore, our framework leverages the expressive and differentiable nature of 3DGS to strike a unique balance between these two conflicting objectives. We introduce a Physical Augmentation Module that injects environmental variations–such as imaging noise, photometric distortions, shadows, and occlusions–into the optimization loop, improving generalization under real-world conditions, addressing Challenge 3. Extensive experiments on both virtual benchmarks and real-world miniature vehicle setups validate our approach: 3DGAA reduces detection mAP from 87.21% to 7.38%, and consistently succeeds across different lighting and viewpoint scenarios, demonstrating its strong physical realism and attack robustness. As shown in Fig. 1, 3DGAA achieves superior physical deployment effectiveness and occupies the optimal region in the realism-robustness space, outperforming existing 2D and 3D baselines.
Our contributions are summarized as follows:
@@
  • We propose 3DGAA, the first adversarial framework to adapt 3D Gaussian Splatting for physical adversarial object generation, enabling joint optimization of geometry and appearance in its native parameter space.
@@
  • We design two novel modules: a physical filtering module to enforce geometric fidelity, and a physical augmentation module to simulate environmental conditions and improve robustness in camera-based object detection systems.
@@
(a) Real-world deployment: object goes undetected
(b) Realism vs. robustness comparison across methods
@@ 2 Related Work
Physical adversarial attacks aim to deceive real-world perception systems by perturbing object appearance in a deployable manner. Early methods [3, 6, 31, 20, 17] mainly adopt 2D patch-based perturbations rendered on printable surfaces. UPC [12] introduces universal physical patches that generalize across scenarios, DAS [33] exploits differentiable mesh rendering [15] to inject adversarial patterns via attention manipulation. Although these approaches offer strong digital attack performance, they often exhibit limited robustness under varying viewpoints, lighting, or physical conditions.
To improve generalization, recent works leverage 3D object representations and render adversarial textures onto full object meshes [37, 35, 33, 32, 26]. In particular, FCA [32] generates full-surface camouflage with physical constraints. Environment-aware enhancements [26, 27] further boost robustness, and recent works such as TT3D [13] explore high-dimensional texture search spaces.
@@
Our approach leverages 3D Gaussian Splatting [16], a differentiable and expressive 3D representation, to jointly optimize both geometry and appearance in a unified 14-dimensional space. This formulation enables physically realistic perturbations that remain effective under variable real-world conditions. Unlike prior works constrained to surface textures, our method introduces adversarial signals through joint optimization of geometry and appearance, enabling physically realistic and adversarially robust object perturbations for evaluating robust perception systems.
@@ 3.1 Preliminaries
Our goal is to generate physically realistic adversarial 3D objects that can degrade the performance of camera-based object detection models under real-world conditions.
@@
The adversarial objective is to perturb the Gaussian parameters such that the object, when rendered into a scene and processed by a detector $D$, minimizes the predicted confidence for its true class. Let $R(\mathcal{G},P)$ denote the rendered image from pose $P$, the core adversarial goal can be formulated as:
@@ 3.2 Overview
As shown in Figure 2, our proposed 3DGAA framework generates adversarial 3D objects through a four-stage pipeline: i 3D Gaussian Generation: Multi-view images and camera parameters are fed into a pretrained backbone to generate an initial 3DGS object. ii Physical Filtering: Structural artifacts and topological noise are removed to enhance geometric plausibility. iii Adversarial Optimization: The filtered object is optimized using an adversarial loss $L_{\text{adv}}$ and a shape consistency loss $L_{\text{shape}}$. iv Physical Augmentation: Differentiable augmentations simulate environmental conditions to improve real-world robustness. Together, these modules produce adversarial objects that are both physically realistic and effective across diverse viewpoints and environments.
@@ 3.3 Physically-Constrained Adversarial Optimization
To dynamically balance attack strength and geometric fidelity, we set $\lambda_{\text{adv}}$ to be inversely proportional to the adversarial loss and $\lambda_{\text{shape}}$ to follow the unnormalized shape deviation. This allows the optimization to prioritize realism when the attack is already strong, and vice versa. Further details are provided in Appendix A.1.
Algorithm 1 summarizes the optimization loop. The 3DGS parameters are iteratively updated via gradient descent, where each iteration involves rendering, applying physical augmentations, computing adversarial and shape losses, and selectively applying gradients to relevant parameter subsets. The dynamic weights $\lambda_{\text{adv}}$ and $\lambda_{\text{shape}}$ are adjusted according to the magnitude of the current loss to balance the strength of the attack and the preservation of the realism.
@@ 3.4 Physical Filtering Module
Despite the flexibility of 3DGS representations, direct optimization over the entire parameter space often results in artifacts such as isolated floating Gaussians, inconsistent scales, and spatial jitter. These issues compromise the visual realism and physical reality of the resulting objects, especially under multi-view rendering. We propose a Physical Filtering Module to refine the optimized 3D object and enforce geometric consistency.
@@
Structural Denoising (SD). We further smooth the object by adjusting the scale of neighboring Gaussians to reduce abrupt local variations. We apply a Gaussian-weighted smoothing filter over scale values, guided by projected image-space distance and opacity. This encourages gradual transitions in apparent size and reduces aliasing artifacts caused by irregular primitive distribution.
@@ 3.5 Physical Augmentation Module
Each transformation is applied in sequence to the rendered image: $\mathcal{T}=\mathcal{T}_{occl}\circ\mathcal{T}_{shadow}\circ\mathcal{T}_{photo}\circ\mathcal{T}_{noise}$. All operations are differentiable and applied online at every training step, allowing gradients to flow through the augmentation pipeline. This enables adversarial optimization to produce samples that remain effective under realistic sensing and environmental distortions. Additional formulation details are provided in Appendix A.3.
@@ 3.6 Selective Dimension Optimization
To support flexible adversarial sample generation under physical constraints, we introduce a manual dimension selection strategy that decouples geometric and appearance parameters in the 3DGS space. Specifically, users can define a subset of dimensions $K\subset\{1,\dots,14\}$ to be optimized, allowing certain parameters, to remain fixed based on deployment requirements. This design enables tailored perturbation control across different real-world scenarios. Implementation details and additional visualizations are deferred to Appendix A.4 for clarity.
@@ 4.1 Experimental Setup
Baseline Methods. We compare 3DGAA against representative 2D and 3D physical attack baselines, including CAMOU[37], UPC[12], DAS[33], FCA[32], DTA[26], ACTIVE[27] and TT3D[13]. For a fair comparison, all methods are adapted to the same synthetic and real-world setups with standard calibration and consistent rendering settings.
@@
Target Models. We evaluate detection attacks on six widely used object detectors: Faster R-CNN [25], Mask R-CNN [10], SSD [19], YOLOv3 [24], YOLOv5 and YOLOv8. For transferable task segmentation, we use DeepLabv3 [5] and FCN [21] with ResNet-50 [22] and ResNet-101 [9] backbones.
@@
Unlike absolute confidence drop, LCR normalizes adversarial effectiveness across different baselines, ensuring a fairer comparison. We also report mean Average Precision (mAP@0.5) to reflect detection performance decline. To quantify physical realism, we adopt image similarity metrics including LPIPS [36], SSIM [34], and PSNR, computed between rendered adversarial objects and their original clean versions.
Implementation Details. All experiments are conducted on a single NVIDIA RTX 4090 GPU. We set the number of epochs in 50, the attack training finished in 1 minute. The two loss terms in our objective–adversarial loss $\mathcal{L}_{\text{adv}}$ and shape consistency loss $\mathcal{L}_{\text{shape}}$–are dynamically weighted during optimization. Specifically, we compute the relative loss magnitudes and normalize them to balance gradients: the weight for $\mathcal{L}_{\text{adv}}$ is inversely proportional to its scaled loss value, while the weight for $\mathcal{L}_{\text{shape}}$ is proportional to its raw value. This dynamic scheme eliminates the need for manual tuning of fixed loss weights.
Table 1: Benchmark Comparison of 3DGAA and Existing baseline Adversarial Attack Methods.
Method None CAMOU[37] UPC[12] DAS[33] FCA[32] DTA[26] ACTIVE[27] TT3D[13] 3DGAA
@@ 4.2 Comprehensive Comparison with Benchmark Physical Attacks
Attack Strength. 3DGAA significantly outperforms all baselines in terms of adversarial effectiveness. It achieves an LCR of 3.5628, far exceeding the second-best method (ACTIVE at 1.3778), and reduces the detector’s mAP from 87.21% to just 7.38%. This indicates that our method not only fools the detector more often, but also leads to lower residual confidence, effectively suppressing recognition. These results validate the effectiveness of simultaneously optimizing both geometric and appearance parameters in the 3DGS space, enabling stronger and more transferable adversarial effects.
Viewpoint Robustness. We evaluate attack consistency by rendering adversarial objects from multiple viewpoints. As shown in Figure 3, the original object is consistently detected (top), while the 3DGAA version evades detection from all views (bottom). We also measure LCR across 12 azimuth angles and 3 distances. The polar heatmap in Figure 4 shows higher success rates at side views and close range, with performance slightly declining under front views or long distances, validating the multi-view training strategy and 3DGAA’s robustness to viewpoint variations.
Together, these results confirm that 3DGAA achieves strong attack success while preserving object realism, and remains robust under diverse camera viewpoints, which are key properties for physically deployable adversarial objects.
Figure 3: Multi-Angle Adversarial Attack Effectiveness: Top row shows original samples with successful detection (green boxes) across different viewpoints. Bottom row demonstrates 3DGAA-generated adversarial textures evading detection (red boxes) at corresponding angles. The consistent performance across viewing perspectives highlights 3DGAA robustness to viewpoint variations.

### v2

@@ Abstract
Camera-based object detection systems play a vital role in autonomous driving, yet they remain vulnerable to adversarial threats in real-world environments. Existing 2D and 3D physical attacks, due to their focus on texture optimization, often struggle to balance physical realism and attack robustness. In this work, we propose 3D Gaussian-based Adversarial Attack (3DGAA), a novel adversarial object generation framework that leverages the full 14-dimensional parameterization of 3D Gaussian Splatting (3DGS) to jointly optimize geometry and appearance in physically realizable ways. Unlike prior works that rely on patches or texture optimization, 3DGAA jointly perturbs both geometric attributes (shape, scale, rotation) and appearance attributes (color, opacity) to produce physically realistic and transferable adversarial objects. We further introduce a physical filtering module that filters outliers to preserve geometric fidelity, and a physical augmentation module that simulates complex physical scenarios to enhance attack generalization under real-world conditions. We evaluate 3DGAA on both virtual benchmarks and physical-world setups using miniature vehicle models. Experimental results show that 3DGAA achieves to reduce the detection mAP from 87.21% to 7.38%, significantly outperforming existing 3D physical attacks. Moreover, our method maintains high transferability across different physical conditions, demonstrating a new state-of-the-art in physically realizable adversarial attacks.
@@ 1 Introduction
Camera-based object detection systems serve as a cornerstone of autonomous driving [2, 1, 37], enabling essential tasks such as obstacle avoidance and lane navigation. However, increasing researches show that these models are highly vulnerable to adversarial examples. In physical world, subtle, structured perturbations often induce misclassifications [10, 4, 11, 30, 46, 28]. This raises serious concerns about the reliability of perception modules in safety-critical scenarios. Thus, systematically exploring such vulnerabilities is essential to understand and improve the real world safety of autonomous driving systems [3, 7, 38, 45, 15, 42, 40, 39, 32, 33, 16, 26].
@@
Existing physical adversarial attacks typically fall into two categories: 2D patch-based methods and 3D texture-level attacks. 2D methods typically apply adversarial patches onto objects [3, 7, 38, 15], achieving high physical realism through minimal appearance changes, but often suffer from poor robustness across viewpoints or physical conditions (Challenge 1). In contrast, recent 3D attacks [45, 42, 40, 39, 32, 33, 26] achieve stronger and more consistent multi-view adversarial performance by directly manipulating surface textures. However, they often rely solely on appearance perturbations, which may result in visual artifacts or unrealistic surface distortions. This limits their applicability in safety-critical settings, where high physical realism is essential for real-world deployment (Challenge 2). This reveals an inherent trade-off between physical realism and adversarial robustness, posing a key challenge for deploying such attacks in real-world scenarios where both are critical (Challenge 3).
@@
To address these challenges, we introduce 3DGAA, an adversarial attack framework that performs joint optimization of geometric and appearance perturbations to generate physically realistic adversarial objects. This is achieved by leveraging 3D Gaussian Splatting (3DGS) [19], a differentiable and compact 3D representation originally designed for photorealistic rendering. Unlike mesh-based or point-based representations, 3DGS encodes both shape and texture in a unified 14-dimensional parameter space, enabling fine-grained and physically consistent adversarial manipulation [14, 35, 34].
@@
3DGAA comprises three key modules that jointly address the trade-off between physical realism and attack robustness: First, a Physically-Constrained Adversarial Optimization stage jointly perturbs geometry and appearance attributes to degrade detector confidence across diverse viewpoints and camera distances. This overcomes the limited perturbation scope of 2D methods and addresses Challenge 1. Second, we propose a Physical Filtering Module that enhances geometric fidelity by removing topological outliers and denoising structural artifacts. This module enforces surface-level consistency and mitigates unrealistic deformations commonly seen in texture-only adversarial models, thereby preserving the visual plausibility of the 3D object from multiple perspectives. It further enhances physical reality, tackling Challenge 2. Furthermore, our framework leverages the expressive and differentiable nature of 3DGS to strike a unique balance between these two conflicting objectives. We introduce a Physical Augmentation Module that injects environmental variations–such as imaging noise, photometric distortions, shadows, and occlusions–into the optimization loop. This improves generalization under real-world conditions, addressing Challenge 3.
Extensive experiments on both virtual benchmarks and real-world miniature vehicle setups validate our approach. Specifically, 3DGAA reduces detection mAP from 87.21% to 7.38%, and consistently succeeds across different lighting and viewpoint scenarios. These results demonstrate its strong physical realism and attack robustness. As shown in Fig. 1, 3DGAA achieves superior physical deployment effectiveness and occupies the optimal region in the realism-robustness space, outperforming existing 2D and 3D baselines. Our contributions are summarized as follows:
@@
  • We propose 3DGAA, the novel adversarial framework to adapt 3D Gaussian Splatting for physical adversarial object generation, enabling joint optimization of geometry and appearance in its native parameter space.
@@
  • We design two novel modules: a physical filtering module leveraging topological pruning and structural denoising to enforce geometric fidelity, and a physical augmentation module to simulate environmental conditions and improve robustness in camera-based object detection systems.
@@
(a) Real-world deployment: object goes undetected.
(b) Realism vs. Robustness comparison across methods.
@@ 2 Related Work
Physical adversarial attacks. In autonomous driving scenarios, physical adversarial attacks aim to deceive real-world perception systems by perturbing object appearance in a deployable manner. Early methods [3, 7, 38, 24, 20] mainly adopt 2D patch-based perturbations rendered on printable surfaces. UPC [15] introduces universal physical patches that generalize across scenarios, DAS [40] exploits differentiable mesh rendering [18] to inject perturbation via attention manipulation. Although these approaches offer strong attack performance, they often exhibit limited robustness under varying viewpoints and lighting.
To improve generalization, recent works leverage 3D object representations and render adversarial textures onto full object meshes [45, 42, 40, 39, 32]. In particular, FCA [39] generates full-surface camouflage with physical constraints. Environment-aware enhancements [32, 33] further boost robustness, and recent works such as TT3D [16] and PGA [26] explore high-dimensional texture search spaces.
@@
Our approach leverages 3D Gaussian Splatting [19], a differentiable and expressive 3D representation, to jointly optimize both geometry and appearance in a unified 14-dimensional space [9, 5, 34, 14, 21]. This formulation enables physically realistic perturbations that remain effective under variable real-world conditions [36, 43]. Unlike prior works constrained to textures, our method introduces perturbation through joint optimization of geometry and appearance, enabling physically realistic and adversarially robust object perturbations for evaluating robust perception systems.
@@ 3.1 Preliminaries
Optimizing target. Our goal is to perturb the Gaussian parameters such that the object, when rendered into a scene and processed by a detector $D$, minimizes the predicted confidence for its true class. Let $R(\mathcal{G},P)$ denote the rendered image from pose $P$, the core adversarial goal can be formulated as:
@@ 3.2 Overview
As shown in Figure 2, our proposed 3DGAA framework generates adversarial 3D objects through a four-stage pipeline: i. 3D Gaussian Generation: Multi-view images and camera parameters are fed into a pretrained backbone to generate an initial 3DGS object. ii. Physical Filtering: Structural artifacts and topological noise are removed to enhance physical realism. iii. Adversarial Optimization: The filtered object is optimized using an adversarial loss $L_{\text{adv}}$ and a shape consistency loss $L_{\text{shape}}$. iv. Physical Augmentation: Differentiable augmentations simulate environmental conditions to improve real-world robustness. Together, these modules produce adversarial objects that are both physically realistic and robust across diverse viewpoints and environments.
@@ 3.3 Physically-Constrained Adversarial Optimization
To dynamically balance attack strength and geometric fidelity, we set $\lambda_{\text{adv}}$ to be inversely proportional to the adversarial loss and $\lambda_{\text{shape}}$ to follow the unnormalized shape deviation. This allows the optimization to prioritize realism when the attack is already strong, and vice versa.
Algorithm 1 summarizes the optimization loop. The 3DGS parameters are iteratively updated via gradient descent, where each iteration involves rendering, applying physical augmentations, computing adversarial and shape losses, and selectively applying gradients to relevant parameter subsets. The dynamic weights $\lambda_{\text{adv}}$ and $\lambda_{\text{shape}}$ are adjusted according to the magnitude of the current loss to balance the strength of the attack and the preservation of the realism. Further details are provided in Appendix A.1.
@@ 3.4 Physical Filtering Module
Despite the flexibility of 3DGS representations, direct optimization over the entire parameter space often results in artifacts such as isolated floating Gaussians and spatial jitter. These issues compromise the physical reality of the resulting objects, especially under multi-view rendering. We propose a Physical Filtering Module to refine the optimized 3D object and enforce geometric consistency.
@@
Structural Denoising (SD). We further smooth the object by adjusting the scale of neighboring Gaussians to reduce abrupt local variations. We apply a Gaussian-weighted smoothing filter over scale dimension, guided by projected image-space distance and opacity. This encourages gradual transitions in apparent size and reduces aliasing artifacts caused by irregular primitive distribution.
@@ 3.5 Physical Augmentation Module
Each transformation is applied in sequence to the rendered image: $\mathcal{T}=\mathcal{T}_{occl}\circ\mathcal{T}_{shadow}\circ\mathcal{T}_{photo}\circ\mathcal{T}_{noise}$. All operations are differentiable and applied online at every training epoch, allowing gradients to flow through the augmentation pipeline. This enables adversarial optimization to produce samples that remain effective under realistic sensing and environmental distortions. Additional formulation details are provided in Appendix A.3.
@@ 3.6 Selective Dimension Optimization
To support flexible adversarial sample generation under physical constraints, we introduce a manual dimension selection strategy that decouples geometric and appearance parameters in the 3DGS space. Specifically, users can define a subset of dimensions $K\subset\{1,\dots,14\}$ to be optimized, allowing certain parameters to remain fixed based on deployment requirements. This design enables tailored perturbation control across different real-world scenarios. For instance, one may constrain geometry when targeting printable objects or preserve opacity to ensure sensor transparency. The strategy also provides practical benefits in training efficiency and transferability, as further validated in Section 4.5. Implementation details and additional visualizations are deferred to Appendix A.4 for clarity.
@@ 4.1 Experimental Setup
Baseline Methods. We compare 3DGAA against representative open-source 2D and 3D physical attack baselines, including CAMOU[45], UPC[15], DAS[40], FCA[39], DTA[32], ACTIVE[33] and TT3D[16]. For a fair comparison, all methods are adapted to the same synthetic and real-world setups with standard calibration and consistent rendering settings.
@@
Target Models. We evaluate detection attacks on six widely used object detectors: Faster R-CNN [31], Mask R-CNN [13], SSD [23], YOLOv3 [29], YOLOv5 and YOLOv8. For transferable task segmentation, we use DeepLabv3 [6] and FCN [25] with ResNet-50 and ResNet-101 [12] backbones.
@@
Unlike absolute confidence drop, LCR normalizes adversarial effectiveness across different baselines, ensuring a fairer comparison. We also report mean Average Precision with 50% threshold (mAP@0.5) to reflect detection performance decline. To quantify physical realism, we adopt image similarity metrics including LPIPS [44], SSIM [41], and PSNR, computed between rendered adversarial objects and their original clean versions.
Implementation Details. All experiments are conducted on a single NVIDIA RTX 4090 GPU. We set the number of epochs in 50, the attack training finished in 1 minute (see Appendix C.3). The two loss terms in our objective–adversarial loss $\mathcal{L}_{\text{adv}}$ and shape consistency loss $\mathcal{L}_{\text{shape}}$–are dynamically weighted during optimization. Specifically, we compute the relative loss magnitudes and normalize them to balance gradients: the weight for $\mathcal{L}_{\text{adv}}$ is inversely proportional to its scaled loss value, while the weight for $\mathcal{L}_{\text{shape}}$ is proportional to its raw value. This dynamic scheme eliminates the need for manual tuning of fixed loss weights.
Table 1: Benchmark Comparison of 3DGAA and existing baseline adversarial attack methods. Our method 3DGAA achieves the best performance across almost all metrics, demonstrating superior physical realism and adversarial robustness.
Method Vanilla CAMOU[45] UPC[15] DAS[40] FCA[39] DTA[32] ACT.[33] TT3D[16] 3DGAA
@@ 4.2 Comprehensive Comparison with Benchmark Physical Attacks
Attack Strength. 3DGAA significantly outperforms all baselines in terms of adversarial robustness in Table 1. It achieves an LCR of 3.5628, far exceeding the second-best method (ACTIVE at 1.3778), and reduces the detector’s mAP from 87.21% to just 7.38%. This indicates that our method not only fools the detector more often, but also leads to lower residual confidence, effectively suppressing recognition. These results validate the effectiveness of simultaneously optimizing both geometric and appearance parameters in the 3DGS space, enabling stronger and more transferable adversarial effects.
Viewpoint Robustness. We evaluate attack consistency by rendering adversarial objects from multiple viewpoints. As shown in Figure 3, the original object is consistently detected (top), while the 3DGAA version evades or misleads detection from all views (bottom). We also measure LCR across 12 azimuth angles and 3 distances. Figure 4 shows higher success rates at side views and close range, with performance slightly declining under front views or long distances, validating the multi-view training strategy and 3DGAA’s robustness to viewpoint variations.
Together, these results confirm that 3DGAA achieves strong attack success while preserving physical realism, and remains effective under diverse viewpoints, which are key properties for physically deployable adversarial objects.
Figure 3: Multi-Angle Adversarial Attack Effectiveness: Top row shows original samples with successful detection (green boxes) across different viewpoints. Bottom row demonstrates 3DGAA-generated adversarial textures evading detection (black boxes) or false detection (red boxes) at corresponding angles. The consistent performance across viewing perspectives highlights 3DGAA robustness to viewpoint variations.